### v2

# CCAT: Magnetic Sensitivity Measurements of Kinetic Inductance Detectors

Abstract
The CCAT Observatory is a ground-based submillimeter to millimeter experiment located on Cerro Chajnantor in the Atacama Desert, at an altitude of 5,600 meters. CCAT features the 6-meter Fred Young Submillimeter Telescope (FYST), which will cover frequency bands from 210 GHz to 850 GHz using its first-generation science instrument, Prime-Cam. The detectors used in Prime-Cam are feedhorn-coupled, lumped-element superconducting microwave kinetic inductance detectors (KIDs). The telescope will perform wide-area surveys at speeds on the order of degrees per second. During telescope operation, the KIDs are exposed to changes in the magnetic field caused by the telescope’s movement through Earth’s magnetic field and internal sources within the telescope. We present and compare measurements of the magnetic sensitivity of three different CCAT KID designs at 100 mK. The measurements are conducted in a dilution refrigerator (DR) with a set of room temperature Helmholtz coils positioned around the DR. We discuss the implications of these results for CCAT field operations.
I Introduction
The Cerro Chajnantor Atacama Telescope (CCAT) [1] is a ground-based submillimeter to millimeter experiment located at an altitude of 5,600 m in the Atacama Desert. CCAT features the 6 m Fred Young Submillimeter Telescope (FYST). Prime-Cam [2], one of FYST’s first-generation science instruments, will observe in the 210–850 GHz range using feedhorn-coupled, lumped-element superconducting microwave kinetic inductance detectors (KIDs). Prime-Cam can host up to seven instrument modules, in two general types. Broadband modules, named for their band center, use broadband, dual-polarization KID arrays. While spectroscopically-enabled modules are designed for imaging spectroscopy, such as the Epoch of Reionization Spectrometer (EoR-Spec) module [3] fielding a Fabry-Perot Interferometer and 210 GHz to 420 GHz non-polarized KID arrays. Each module contains three detector arrays, designed for that module’s scientific goal. The KID arrays are fabricated on 550 m silicon-on-insulator wafers and operated at around 100 mK in individual instrument modules. We aim to deploy Prime-Cam with two broadband instrument modules centered at 280 [4, 5, 6, 7] and 350 GHz [8] in 2026, followed by an additional broadband module at 850 GHz [9] and the EoR-Spec module.
During observations, the telescope will conduct wide-area surveys (5–20,000 deg²) at scan speeds of degrees per second. The survey area and scan strategy depends on the science targets and are being finalized at the time of writing. In these observations, the detectors are exposed to magnetic field variations from the telescope’s motion through Earth’s magnetic field and from internal sources within the telescope. Previous studies have shown that magnetic fields impact superconducting materials which will affect KID performance [10, 11, 12, 13, 14]. Here, we investigate the effect of an external DC magnetic field on CCAT KIDs.
II Experimental Set-up
For the magnetic sensitivity measurements, we used witness chips fabricated alongside the CCAT 280 GHz [4] and low-band EoR-Spec [3] detector arrays with the same lithographic processes and materials. Three witness chips were packaged for this study: two 280 GHz chips (15 mm × 8 mm) made from aluminum (Al) and titanium-nitride (TiN), and one Al EoR-Spec chip (15 mm × 15 mm). The Al and TiN 280 GHz pixels (shown in the left and center panels of Figure 1) are lumped element polarimeter designs where one pixel has two absorbers measuring two orthogonal linear polarizations. Each detector is capacitively coupled to the transmission line by a coupling interdigitated capacitor (IDC). The resonance frequency () is set by an inductor (with inductance L) which is also the absorber and parallel IDC (with capacitance C) as . The large difference in the kinetic inductance per square of the two materials led to a much finer linewidth and a meandered absorber for the Al detector, while the TiN absorber [15] incorporated short patches of Al to tune detector responsivity[4]. While also having a lumped element design, the EoR-Spec detectors (shown in the right panel of Figure 1) do not require polarization sensitivity and use a meandered Hilbert curve absorber[16]. Each chip was mounted in a separate copper test box with rubber cement, and read out via gold-plated brass SMA or SMP connectors and Al wire bonds. The box was sealed with a copper lid and brass screws. Figure 2 shows the EoR-Spec chip integrated into the base of the copper test box, as an example.
The cryogenic testing was performed in a Bluefors LD dilution refrigerator (DR). All three witness chips were tested together, with their copper test boxes mounted on the DR mixing chamber (MXC) plate in close proximity. The TiN 280 GHz and EoR-Spec boxes were stacked near the center on the top of the mixing chamber plate, while the Al 280 GHz box was mounted near the center on the bottom of the mixing chamber plate. This arrangement placed all boxes close to the center of the applied magnetic field and near each other to ensure they experienced the same field amplitude. In two DR cooldowns, the Al 280 GHz box was flipped upside down to confirm that the measurements were independent of the chip’s upside-down orientation. There are three independent cold readout channels in the DR, with each chip read out through its own channel. Each channel is similar to that planned for Prime-Cam [17], with cryogenic low-noise amplifiers at the 4 K stage on the RF output chain, various cryogenic attenuators on the RF input chain, and cryogenic coaxial cables. A custom-packaged Radio Frequency System-on-Chip (RFSoC) [18, 19] serves as the main room-temperature readout electronics, and an Agilent E5071C vector network analyzer was used to assist the measurements.
A set of Helmholtz coils was mounted on the DR frame outside the cryostat to provide an adjustable, uniform magnetic field as shown in Fig 3. This set of coils was previously used for similar measurements on other devices [20, 21]. A Lakeshore 460 three-channel gaussmeter was used to calibrate the coil performance by comparing calculated field strength at various drive voltages with measured values, and to confirm field uniformity at the center of the coil set. Additional measurements confirmed that the DR shells at various temperature stages do not provide additional magnetic shielding.
A room-temperature mu-metal magnetic shield was installed outside the DR and removed before the magnetic measurements. Following each set of measurements, the magnetic shield was reinstalled while the detectors were warmed up above the critical temperature, , of the superconducting aluminum (between 1.2 K and 1.4 K), before cooling back down. The importance of a shielded magnetic environment during cooldown past the KID’s critical temperature has been well documented in previous publications [13, 10] and is also discussed in Section III. After cooling to base temperature, the DR mixing chamber was servoed to 100 mK. The resonators from the test chips were individually tuned with different input RF powers to achieve high sensitivity while avoiding bifurcation [22]. The following analysis uses all of the available detectors on each chip which amounts to five Al 280 GHz detectors, two TiN 280 GHz detectors, and five low-band EoR-Spec detectors.
III Measurements and Analysis
To measure the magnetic sensitivity of the KIDs, the voltage across the coils was ramped up from - V (- A) in steps of V, this translated to a range of approximately - T in steps of T. At each external magnetic field magnitude, , we waited for seconds and then commenced a set of five scattering forward transmission () sweeps centered at each resonator’s nominal resonance frequency, .
Figure 4 shows an example of the measured for a sample Al 280 GHz detector from two consecutive sweeps, where in each sweep the magnetic field was ramped up in magnitude and then back down. To better understand the detectors’ response to the external magnetic field, we extract the resonant frequency, , and internal quality factor, , by fitting the data to the following resonator model [23],
Above, the contributions from the readout network and cable are captured in , , the cable delay (), and phase offset (), while the resonator parameters , , and , are the total quality factor, external coupling quality factor, and resonant frequency, respectively. The internal quality factor of the resonator is obtained from and by the following relation,
is the coupling quality factor, which is purely real and can be determined from which is complex. Following an initial fit with the full set of frequency data to fix the cable components of the model, a much narrower window of data was then fit to a full nonlinear resonator model as described in [22] and shown below.
In this equation, is the resonator detuning in linewidths accounting for the non-linear kinetic inductance given by the solutions to the equation
Here, depicts the degree of non-linearity in the resonator. Since we took multiple target sweeps at each magnetic field, the values presented are the medians, and the error bars shown in the plots are defined as the standard deviations. With the gaussmeter, we estimated the precision of the applied field to be about 5 µT. For clarity, the uncertainty on the magnetic field strength are not shown in any of these plots.
From these fits we obtain two main observables, the fractional change in resonant frequency, and the resonator’s internal quality factor. We first investigate the fractional frequency shift via,
Above, is the zero field resonant frequency. This comes from a set of target sweeps performed at T, and while the cryostat was still enclosed by -metal after reaching 100 mK for the first time. Plotted in Figure 5 is versus over both vertical directions (positive with the field pointing vertically upward and negative downward). It is evident in Figure 5 that there is a hysteresis effect in which the resonator properties change differently when the magnetic field strength is increased compared to when it is decreased. This is further demonstrated in the quality factor of the resonator. In Figure 4, the width and depth of the resonator changes as a function of magnetic field, and on subsequent sweeps the dip depth does not fully recover to its initial value of dB.
We investigate these effects in more detail by looking at an individual resonator that is characteristic of the whole set. Fractional shifts in and are plotted for an EoR-Spec detector in Figure 6. A similar hysteretic degradation of the quality factor is seen. We hypothesize that this is due to the formation of trapped magnetic flux or magnetic vortices within the superconducting films as they are exposed to these magnetic fields, introducing excess loss and decreasing the amount of stored energy available; however, further investigation is needed to confirm this.
We also study the effects of a magnetic field parallel to the detector absorber plane by placing the coils in a horizontal orientation. In the same manner as the other data-set we fit a non-linear resonator model to the data and extract the fractional shift in resonant frequency. As shown in Figure 7, the effects of the horizontal field on are a few orders of magnitude lower than that of the vertical orientation.
As an aside, we performed a set of measurements without the magnetic shield in place while cooling down below the detectors’ . We found that there is a bump in the when sweeping through the negative field, where the field direction points vertically downward as shown in Figure 8. We hypothesize that this effect is caused by trapped flux induced by the Earth’s magnetic field as the temperature passes through and the chips become superconducting.
IV Discussion
We find that, when subjected to external magnetic fields, the KIDs exhibit degraded internal quality factors and reduced responsivity, with the effect dominated by magnetic fields perpendicular to the detector plane. We now discuss the implications this has on CCAT observations.
First, we note that we were unable to fully decouple the hysteresis observed in the chip from hysteresis originating in the DR. Although the chip packaging was constructed from non-magnetic materials, components in the DR made of stainless steel, which are not feasible to replace, can still introduce magnetic hysteresis. Future studies of magnetic field effects on KIDs would benefit from either a cold Helmholtz coil that couples only to the test chip or a cryogenic magnetometer to monitor the local field around the test device.
At the observing site, the maximum expected magnitude of the Earth’s magnetic field is approximately 25 T111Estimated using the NOAA magnetic field calculator at https://www.ngdc.noaa.gov/geomag/calculators/magcalc.shtml at the FYST location (22∘59′09′′S, 67∘44′25′′W).. Changes in the magnetic field strength that the detectors experience due to the telescope moving through the Earth’s field should be less than twice this value. Taking the 50 T as an upper bound to the possible change in field would yield a corresponding fractional frequency shift of less than , as shown in Figure 7. Using the previously reported detector sensitivity for the Al 280 GHz detectors [5], under an optical loading of approximately 5.6 pW in the best weather quartile, this fractional frequency shift corresponds to an optical signal of about 0.3 pW, which can be considered an extreme upper bound.
Inside the cryostat, a Cryoperm A4K shield222Amuneal Mfg. Corp., Philadelphia, PA 19124 covers the back of the instrument modules and provides magnetic shielding at 4 K [24]. Although the shielding factor has not been directly measured, simulations estimate it to be greater than 200, consistent with other experiments employing similar magnetic shielding schemes [25, 26]. Additionally, any shifts due to the magnitude field ought to occur as a long baseline common mode signal across all detectors in the module, meaning they will appear in the data in a manner similar to the atmospheric loading. Considering the relatively small changes in magnetic field magnitude, the module’s magnetic shielding, and the common-mode nature of the signal, we conclude that the magnetic field variations encountered while scanning through Earth’s field during telescope operations are expected to have a negligible impact on the detectors.
Acknowledgments
The CCAT project, FYST and Prime-Cam instrument have been supported by generous contributions from the Fred M. Young, Jr. Charitable Trust, Cornell University, and the Canada Foundation for Innovation and the Provinces of Ontario, Alberta, and British Columbia. The construction of the FYST telescope was supported by the Großgeräte-program of the German Science Foundation (Deutsche Forschungsgemeinschaft, DFG) under grant INST 216/733-1 FUGG, as well as funding from Universität zu Köln, Universität Bonn and the Max Planck Institut für Astrophysik, Garching. The construction of EoR-Spec is supported by NSF grant AST-2009767. The construction of the 350 GHz instrument module for Prime-Cam is supported by NSF grant AST-2117631. S. Walker acknowledges support from the National Science Foundation under Award No. 2503181. S. Chapman acknowledges both the Natural Sciences and Engineering Research Council of Canada and the Canadian Foundation for Inovation.